OpenStudy (daisy.xoxo619):

Find a point-slope form equation for the line through the point (-4, 3) with a slope of -2/3.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ -4}}\quad ,&{\color{blue}{ 3}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies-\cfrac{2}{3} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (daisy.xoxo619):

y-3=-2/3(x+4) is that correct ?

OpenStudy (jdoe0001):

yeap, that's correct :)